. A circle of radius 2 cm is cut out from a square piece of an aluminium sheet of side 6 cm. What is the area of the left over aluminium sheet? (Take π = 3.14)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question it is given that,
Radius of circle = 2 cm
Square sheet side = 6 cm
First we have to find out the area of square aluminium sheet and circle to find out the remaining area.
Now,
Area of the square = side^2
= 6^2
= 36 cm^2
Area of the circle =πr^2
= 3.14 × 2^2
= 3.14 × 2 × 2
= 3.14 × 4
= 12.56 cm2
Now,
The area of the remaining part = Area of aluminum square sheet–area of circle
= 36–12.56
= 23.44 cm^2

Thanks